in a right angled triangle abc, b is right angle, side ab is half of the hypotenuse. ae is parallel to median bd and ce is parallel to ba. what is the ratio of length of bc to that of ae?

Grade:11

1 Answers

Ashutosh nayak
7 years ago
See ac=h,ab=h/2so ac=3½/2 so from trigonometry knowledge angleDBC=30°.then draw a perpendicular to CE .let point be F .so angleEAF=angle DBC.=30°thenAF=BC=3½/2so find AEasAE=AFsecø.sobc/ae=(3½/2)h/h=3½..